# End-to-End 4D Heart Mesh Recovery Across Full-Stack and Sparse Cardiac MRI

### v1

@@ 4.2 Motion Estimation from a Few Slices
In this section we present results under the online few-slice scenario that has been the main motivation for this work. In Fig. 2, we show two representative motion reconstruction results using only the central slice as input: one from a normal (NOR) heart and another from a subject with dilated cardiomyopathy (DCM). Because of its stronger contractile function, the normal heart exhibits larger deformations than the DCM heart, making motion prediction more challenging, which is reflected by larger magnitudes in the error maps. Nevertheless, our method delivers accurate predictions in both cases. Furthermore, the segmentation results on slices at the apex and base location demonstrate that our model can provide reliable deformation estimation even in regions far from the slices used for motion inference. The complete reconstructed sequences together with the corresponding volume–time curves are shown in Fig. 3.
We report quantitative results on the whole test set in Tabs. 1 and at the top of Tab. 2. The first contains chamfer distances and dice scores and the second cardiac function index predictions. For consistency, we generated all these results using the central 1 or 5 slices even though slices at any position could be used, as shown in Section 4.4. TetHeart consistently outperforms our four baselines without requiring ground-truth segmentations or contours as input. Instead, it infers cardiac motion directly from raw CMR images without any time-consuming test-time optimization. As a result, it delivers fast inference—12 FPS on a single V100—with a potential for further acceleration by adopting a lightweight encoder architecture and leveraging model quantization techniques to meet the real-time constraints. Even when using a single slice, TetHeart delivers meaningful indices. All of this points toward its suitability for eventual clinical deployment.
Although we incorporated into Ours-Mesh and Ours-SDF the same mechanisms for handling sparse features as in our complete method, the latter still delivers better results. This confirms that the tetrahedral representation we use is more effective at preserving spatial information within the object, making it a better choice for inferring 3D motion when only limited observations are available. To quantify this, we plot in Fig. 4 the performance of all three methods as a function of the number of slices used. The fewer, the greater the gap between our complete approach and the other two. Nevertheless, both variants still surpass prior approaches by a clear margin, demonstrating that design components such as the AFA module and our training strategy form a general and effective solution for handling incomplete slice scenarios.
The comparatively poor performance of 4DMR can be attributed to its reliance on a simple MLP architecture, which may lack the representational power to generalize well to the large and heterogeneous datasets considered in this study. Moreover, its test-time optimization strategy does not generalize well to severely under-sampled inputs. The weaker performance of MR-Net likely stems from the fact that it reconstructs meshes solely from contour points, thereby neglecting the rich information available in cine images.
4.3 Motion Estimation from Full Stacks
Figure 5: Full-Slice Comparison. (Left) We predict 3D meshes by deforming from ED to ES frame. Color indicates the magnitude of point-to-surface error in $m$. (Right) Segmentation results on slices at the apex and base location. Note 4DMR and DeepMesh can only output for myocardium.
Since we had to modify existing techniques to compare them to ours in the few-slice scenario, we now compare our method to the original baselines on the full set of slices they were designed to operate on and report the results in Tab. 3 and the bottom of Tab. 2. In Fig. 3 and 5, we show qualitative results on the same hearts as in Fig. 2, but now using all slices as input instead of only one.
Remarkably, even though the main driver of this work was operating on as few slices as possible, we also consistently outperform the baselines in the full-slice scenario. We attribute this to several factors. Consistent with the previous analysis, the relatively poor performance of MR-Net may be attributed to its neglect of the rich information available in cine images. Using predicted contours resulted in slightly inferior performance, indicating that such a two-stage automated pipeline is prone to error accumulation, a finding consistent with the conclusions in the original paper [26]. MulViMotion estimates voxel-wise motion fields and reconstructs meshes from warped segmentation maps, which is inherently less accurate than directly modeling meshes. Compared to DeepMesh, our model benefits from a shared encoding network, allowing the deformation branch to leverage informative features from the reconstruction branch. Moreover, our AFA module dynamically aggregates spatially adjacent features, further enhancing prediction quality.
4.4 External Evaluation: To Generalize without Retraining
We now test the generalizability of our model to the CHUV and 4DM datasets without retraining.
In Fig. 6, we show sequence predictions on CHUV with the spatial position of the real-time captured 2D slice visualized in the left. As can be seen, TetHeart reconstructs a plausible heart motion across the cardiac cycle under both rest and exercise situation. Moreover, the predicted volume–time curves further confirm the physiological plausibility of our motion reconstruction. In the exercise condition, the heart exhibits a shorter ES–ED recovery period and a reduced duration of each cardiac cycle, which aligns well with expectation.
As 4DM provides full-sequence annotations, but only for the myocardium. we evaluate the full-sequence prediction performance for the myocardium subclass, rather than focusing solely on the ES frame and report the results in Tab. 4. Here, check marks denote the dataset used for training. Oracle refers to training and testing directly on 4DM, serving as an upper bound for performance.
As shown in the table, our model demonstrates the best generalization abilities, achieving top performance in both few-slice and full-slice settings. In contrast, 4DMR exhibits poor generalization, likely due to significant distribution shifts across datasets and the limited representational capacity of its simple MLP architecture. While MulViMotion and DeepMesh show some generalization ability, their lower baseline performance causes them to lag behind our method.
Moreover, training with a larger and more diverse dataset improves generalization. And consistent with the findings from previous evaluations, representing the heart with tetrahedra enables more effective spatial information retention and exchange, particularly in scenarios with limited slice input.
@@ 4.4 External Evaluation: To Generalize without Retraining
Figure 6: Motion sequence prediction on the CHUV dataset. (Left) Spatial configuration of the initial SAX stack and the online observations. (Middle) Cardiac motion sequences for each subject under rest and exercise conditions. (Right) Volume–time curves.
4.5 Ablation Studies
Finally, we conduct ablation studies to evaluate our design choices. First, we examine the role of the AFA module. Simply stacking 2D features achieves comparable performance to AFA in the full-slice setting, but it fails to generalize to scenarios with variable numbers and positions of 2D slices. This highlights the necessity of a more flexible mechanism like AFA. Next, we evaluate the importance of spatial information in the AFA module. If spatial position encodings are removed from the attention computation, the model’s performance degrades moderately, indicating that spatial information is crucial for accurate feature aggregation. However, if we use only position encodings as the query and discard image features entirely, the performance drops substantially. This shows that both image features and spatial coordinates are essential and must work together for effective motion prediction. When training the motion branch, if we do not share parameters with the reconstruction branch and instead initialize the deformation branch randomly, the model performance drops significantly. This demonstrates the effectiveness of our parameter-sharing strategy: the 3D knowledge learned during reconstruction is successfully transferred to motion modeling.
Lastly, we ablate the impact of the distillation loss. Without it, the model achieves slightly better results in the full-slice setting, but the features encoded from sparse 2D inputs are less discriminative, leading to poorer performance under few-slice conditions. When the distillation loss is combined with our aggressive sampling strategy, the model achieves consistently stronger results across all settings, demonstrating the importance of this training scheme for improving robustness and generalization.
Table 5: Ablation study results on M&Ms dataset.
  Method M&Ms Myocardium CD (mm2) $\downarrow$
@@ 4.5 Ablation Studies
Ours 9.76 7.93 4.11
- AFA module - - 4.22
- position embedding 11.23 9.04 4.67
- image feature in query $Q$ 16.67 12.37 6.44
- shared encoding network 20.66 11.63 6.43
- distillation loss 10.34 8.12 4.03
5 Conclusion
We have proposed TetHeart, a unified framework for reconstructing heart motion from either full stacks of 2D MRI slices or from sparse sets, down to a single slice. We demonstrate, for the first time, that full 3D heart shape and motion can be inferred in sparse real-time CMR slices, making our approach suitable for guidance of interventional procedures. This breakthrough shifts cardiac reconstruction from retrospective analysis to live clinical use, laying the foundation for real-time digital twins of the beating heart. This capability not only has the potential to enable precise, image-guided interventions today, but also points to a future where digital twins of organs in motion guide therapies across medicine—adaptive, predictive, and personalized at the point of care.
[…]

### v2

@@ Appendix C Baselines
Table S1: Zero-shot evaluation results on public 4DM dataset. Check marks denote the dataset used for training. Oracle is trained and evaluated on 4DM.
Appendix D Complementary Experiment Results
High-Res Verision of Fig. 4. In Fig. S1, we provide a higher resolution version of Fig. 4 for better view.
Disease Definition. A normal heart should satisfy the following criteria: LV EF $50\%$ and RV EF $40\%$. A subject with dilated cardiomyopathy should have LV EF $40\%$. This is not a strict definition; a more rigorous classification criterion would require a comprehensive analysis of the ratio between volume and body surface area, along with other indicators. However, verifying whether the predicted ejection fraction from TetHeart satisfies this approximate criterion can further demonstrate the practicality of our model.
As could be seen from Fig. S1, even with a single slice, TetHeart could predict cardiac function indexes that meet the criteria. This highlights our model’s practicality.
Visualization Results Using Full Slices. In Fig. S2, we show qualitative results on the same hearts as in Fig. 3. But now using full slices rather than using a single slice.
Appendix E External Evaluation Results
iCMR Dataset. We provide a higher resolution version of Fig. 6 in Fig. S3 for better view.
@@ Appendix E External Evaluation Results
Figure S3: Motion sequence prediction on the iCMR dataset. (Left) Spatial configuration of the initial SAX stack and the online observations. (Middle) Cardiac motion sequences for each subject under rest and exercise conditions. (Right) Volume–time curves.
4DM Dataset. As 4DM provides full-sequence annotations, but only for the myocardium. we evaluate the full-sequence prediction performance for the myocardium subclass, rather than focusing solely on the ES frame and report the results in Tab. S1. Here, check marks denote the dataset used for training. Oracle refers to training and testing directly on 4DM, serving as an upper bound for performance.
As shown in the table, our model demonstrates the best generalization abilities, achieving top performance in both few-slice and full-slice settings. In contrast, 4DMR exhibits poor generalization, likely due to significant distribution shifts across datasets and the limited representational capacity of its simple MLP architecture. While MulViMotion and DeepMesh show some generalization ability, their lower baseline performance causes them to lag behind us.
Moreover, training with a larger and more diverse dataset improves generalization. And consistent with the findings from previous evaluations, representing the heart as tetrahedra enables more effective spatial information retention and exchange, particularly in scenarios with limited slice input.
Appendix F Extended Ablation Studies
The Impact of AFA Configuration. In Tab. S2, we show the impact of the number of slices in the AFA module and the number of selected positions per slice on both performance and speed (measured on an NVIDIA V100 GPU). As shown in the table, our default configuration achieves a good balance between accuracy and efficiency. Further increasing the number of slices or selected positions brings only limited performance gains but leads to a slowdown. One thing worth mentioning is that to implement our AFA module we use a relatively simple strategy of sorting and selecting the top-k elements. Therefore, for $|S|=1,|G|=5^{2}$ and $|S|=3,|G|=3^{2}$, although the two configurations use a similar number of features, the latter runs slower because it performs sorting multiple times. More advanced strategies or precomputation could potentially narrow this gap for further acceleration.
  Model $|S|$ $|G|$ Myocardium CD & FPS ($mm^{2}$ $\downarrow$ & $img/s$ $\uparrow$)
@@ Appendix F Extended Ablation Studies
  TetHeart 1 $1^{2}$ 10.34 15.1 8.45 15.1 4.56 15.1
1 $3^{2}$ 9.98 15.0 8.22 14.9 4.43 14.9
1 $5^{2}$ 9.93 14.8 8.13 14.7 4.39 14.6
3 $1^{2}$ 10.32 15.0 8.19 13.0 4.37 13.0
3* $3^{2}*$ 9.76 14.9 7.93 12.1 4.11 12.0
3 $5^{2}$ 9.73 14.7 7.86 10.8 4.05 10.7
5 $1^{2}$ 10.27 15.0 8.17 11.9 4.33 11.8
5 $3^{2}$ 9.74 14.8 7.81 10.4 4.09 10.4
5 $5^{2}$ 9.72 14.7 7.83 8.7 4.10 8.7
7 $5^{2}$ 9.76 14.6 7.85 8.6 4.07 7.3
Table S2: Impact of different AFA configuration evaluated on the M&Ms dataset. $|S|$: the number of slice. $|G|$: the number of positions selected from each slice. *: default setting.
The Influence of the Input Slice Position. In the main text, for ease of comparison and to ensure fairness, we use central slices as model inputs to evaluate the performance of different methods. Here we investigate how the input slice position affects the accuracy of motion inference. The results are shown in Tab. S3. The numbers indicate the offset of the input slice relative to the central slice: - denotes a shift toward the apex, and ++ denotes a shift toward the base. For example, suppose we have total 11 slices, then the central slice is the 6-th slice, the +1+1 slice is the 7-th slice, the 1-1 slice is the 5-th slice.
As seen from the table, although the model still achieves good motion inference performance, its accuracy steadily decreases as the offset increases. These experimental results suggest that slices at different positions contain different amounts of motion information, which may be more focused on local regions. If only single slice is available, selecting it close to the central position yields better overall motion inference. This conclusion provides guidance on choosing suitable scan positions in real-world intervention workflows. For multiple slices, we could conduct similar experiments to explore which slice combinations are most effective. However, due to the large number of possible combinations, we provide only one general conclusion here: using an appropriate slice sampling interval can improve model performance compared with using adjacent slices.
  Slice Position -3 -2 -1 Central +1 +2 +3
  Myo CD ($mm^{2}$ $\downarrow$) 15.76 13.62 10.01 9.76 10.29 10.81 12.22
Table S3: Impact of input slice position evaluated on the M&Ms dataset. The numbers indicate the offset of the input slice relative to the central slice. - denotes a shift towards the apex, and ++ denotes a shift towards the base.
Category-Wise Ablation Results. In Tab. S4, we present more detailed results on the unified dataset as a supplement to Tab. 4 in the main text. As shown in the table, across all datasets and all categories, the conclusions remain consistent with those in the main text: the design of the AFA module, parameter sharing, and the use of distillation loss all contribute to improving the final performance of our model.
  Method Myocardium CD (mm2) $\downarrow$ LV CD (mm2) $\downarrow$ RV CD (mm2) $\downarrow$
1-Slice 5-Slice Full-Slice 1-Slice 5-Slice Full-Slice 1-Slice 5-Slice Full-Slice
ACDC
Ours 15.24 10.22 6.63 23.65 18.31 11.51 40.64 32.61 22.15
- AFA module - - 6.55 - - 11.58 - - 22.30
- position embedding 16.60 11.33 6.84 25.32 20.14 12.73 42.23 34.19 23.59
- image feature in query $Q$ 24.10 15.96 9.39 36.76 28.39 17.44 58.31 43.21 32.32
- shared encoding network 32.23 15.24 10.01 49.12 28.99 18.59 75.92 44.81 34.44
- distillation loss 16.13 10.48 6.49 25.07 18.78 11.28 43.08 33.43 21.71
M&Ms
Ours 9.76 7.93 4.11 12.37 9.72 4.85 21.42 17.39 13.68
- AFA module - - 4.22 - - 4.79 - - 13.74
- position embedding 11.23 9.04 4.67 13.35 10.34 5.67 22.72 18.84 14.93
- image feature in query $Q$ 16.67 12.37 6.44 19.76 14.17 7.82 33.63 23.81 20.60
- shared encoding network 20.66 11.63 6.43 24.56 14.24 7.77 41.80 24.12 20.45
- distillation loss 10.34 8.12 4.03 13.11 9.94 4.75 22.70 17.79 13.41
M&Ms-2
Ours 10.12 7.30 4.02 13.89 9.96 5.04 20.27 15.58 10.82
- AFA module - - 4.17 - - 5.07 - - 10.76
- position embedding 11.93 8.41 4.74 14.98 10.40 5.69 21.91 16.72 11.65
- image feature in query $Q$ 17.34 11.38 6.31 22.32 13.57 7.57 32.65 21.40 15.49
- shared encoding network 21.49 11.09 6.54 26.96 13.64 7.85 39.44 20.73 16.07
- distillation loss 10.95 7.47 4.11 15.01 10.19 5.05 21.89 15.94 11.06
Table S4: More comprehensive ablation study results on the unified dataset.
Appendix G Limitation
Our method achieve promising results on both full-stack and few-slice settings, surpassing previous methods. However, by examining our own model, we can observe that although we have proposed several techniques, there still remains a performance gap between motion reconstruction using only a few slices versus the full stack. We plan to introduce physical constraints as prior knowledge into the model to reduce this gap. Another choice is to use a generative model to generate the mesh sequence progressively to complement our current feed-forward reconstruction process. However we need to strike a good balance between the reconstruction speed and quality as we are targeting real-time applications. Although we report quantitative results for the few-slice setting on the Unified dataset and achieve the best performance, we can only provide qualitative dynamic reconstruction results on the iCMR dataset because the ground-truth cardiac shapes for single slices are not available. Identifying an effective evaluation metric for assessing the quality of reconstructed mesh sequences in this setting would be very helpful for future clinical deployment. These are all interesting research directions, and we leave them for future work.